# Meta4XNLI: A Crosslingual Parallel Corpus for Metaphor Detection and Interpretation

### v1

@@ 3.3 Resulting Dataset
The parallel data for this task comprises a total of 13320 sentences annotated at token level, since we approached metaphor detection as a sequence labeling task, following the criteria of the cited previous work in Section 2.
With respect to Spanish annotations, there is a total of 1155 metaphorical tokens in premises and 1139 in hypotheses. Out of the 13320 sentences, 1873 contain at least one metaphorical expression, which constitutes the 14% of the whole dataset. This information is detailed in Table 4 according to the source datasets partitions. A higher proportion of metaphors in premises can be noticed. This might be due to the fact that premises are of larger length than hypotheses, as indicated in Conneau et al. (2018). In addition, premises were collected from existing utterances, while hypotheses were generated from crowd-source workers in response to a given set of premises. Therefore, hypotheses sentences might tend to present shorter length and lower complexity.
@@
Regarding English annotations, we can observe a similar trend to that of ES in Table 5. A total of 3330 tokens were labeled as metaphor and 2736 sentences contain at least one metaphorical instance out of the total 13320 sentences. Premises show a higher metaphor ratio than hypotheses as in ES annotations. Additionally, we observe a larger amount of labeled metaphors in EN, caused by the different annotations processes specified in Subsection 3.2. Since VUAM contains a significantly higher amount of labeled metaphorical expressions, the MLM fine-tuned with this dataset predicted many ambiguous metaphors, which we subsequently removed in manual revision. These discrepancies are not only noticeable in annotations but also in experiments results. Moreover, this gives an account of how guidelines for metaphor identification labeling are open to discussion and clarification, due to the subjective and nuanced nature of this cognitive-linguistic phenomenon.
@@
Annotations for this task were developed at premise-hypothesis level. As shown in Table 6, the average percentage of pairs with metaphors relevant to the inference relationship is 12%. This figure remains steady throughout each source dataset and inference labels. esXNLI shows a higher number of metaphor occurrence that might be caused by the difference of text domains and sentence characteristics with respect to XNLI data. Regarding non-relevant cases, we do not exploit them in the experiments in order to be able to analyse more clearly whether metaphor presence impacts models performance. We keep the same sample distribution for both languages to develop the experiments.
@@ 3.4 Inter-annotator Agreement
We selected a subset of 1000 pairs to the assess annotation quality in ES. The subset was labeled by another Spanish native speaker with a linguistics background. This annotator labeled the dataset subset from scratch and according to MIPVU guidelines in the case of detection, and the annotation process specified in Subsection 3.2 for interpretation. We computed Cohen’s Kappa score Cohen (1960) and obtained 0.74 in premises and 0.77 in hypotheses sentences for detection; for interpretation via the NLI task the obtained score was 0.64. Thus, these scores show a substantial agreement and are comparable to those obtained in the annotation of other datasets for metaphor detection and/or interpretation Steen et al. (2010); Sanchez-Bayona and Agerri (2022); Maudslay and Teufel (2022); Badathala et al. (2023) thereby demonstrating the consistency of our annotations. Still, annotating metaphorical language remains a rather subjective task that requires exhaustive work and precise guidelines.
4 Experiments
In this section we present the experimental setup designed with the objective of testing the capabilities of multilingual MLMs on metaphor detection in cross-domain, crosslingual and multilingual settings. Furthermore, we also experiment with their ability to perform NLI when the correct inference requires understanding of metaphorical language.
4.1 Experimental setup
With the aim of developing cross-lingual experiments, we chose the multilingual language models and checkpoints that obtained best results in Sanchez-Bayona and Agerri (2022): mDeBERTa (base) and XLM-RoBERTa (large) Conneau et al. (2020), the multilingual versions of DeBERTa He, Gao, and Chen (2021) and RoBERTa Liu et al. (2019) respectively.
@@ 4.1 Experimental setup
Taking advantage of previous available resources and the corpus we present in this work, we conducted a series of experiments to evaluate and fine-tune MLMs. The configuration of cross-domain experiments is specified in Sanchez-Bayona and Agerri (2022). For the other three setups, monolingual, multilingual and zero-shot cross-lingual, we performed hyperparameter tuning for batch size (8, 16, 36), linear decay (0.1, 0.01), learning rate (in the [1e-5-5e-5] interval), sequence length of 128 and epochs from 4 to 10. A warm-up of 6% is specified. The results of the hyperparameter tuning showed that after 4 epochs development loss started to increase, so results reported here are obtained with 4 epochs, batch size of 8, weight decay of 0.1 and learning rate of 5e-5.
@@
Cross-domain: the aim of this set of experiments is to evaluate the performance of MLMs fine-tuned with CoMeta and VUAM datasets on Meta4XNLIES and Meta4XNLIEN, respectively, since each dataset contains texts from different domains. The motivation is to explore the impact of text features and genres on the performance, as well as annotation criteria Aghazadeh, Fayyaz, and Yaghoobzadeh (2022); Lai, Toral, and Nissim (2023). To do so, we chose the models with best performance from monolingual experiments developed in Sanchez-Bayona and Agerri (2022). We conducted the evaluation on various data splits: within each source dataset, XNLIdev, XNLItest and esXNLI, we evaluated premises and hypotheses separated and combined. This is due to the dissimilarities and the unequal distribution of metaphorical expressions between premises and hypotheses sentences mentioned in Subsection 3.3.
@@
Monolingual: this scenario comprises the fine-tuning and evaluation of the MLM on Meta4XNLIES and Meta4XNLIEN separately. To accomplish this task, we split Meta4XNLI into train, development and test sets (0.6-0.2-0.2). We equally distributed the data to ensure each partition is balanced in terms of source datasets sentences and metaphor occurrence. Data statistics are detailed in Appendix A. These partitions will be used as well in subsequent multilingual and zero-shot cross-lingual scenarios. In addition to fine-tuning and evaluation on Meta4XNLIES and Meta4XNLIEN, we evaluated each trained monolingual model with the test sets of CoMeta and VUAM, following the same reasoning as in cross-domain experiments.
@@
Multilingual: the purpose of these experiments is to explore whether MLMs benefit from being trained on data in multiple languages. In this case, we combined Meta4XNLIES and Meta4XNLIEN train splits to fine-tune the models. Subsequently, we evaluated the trained models on each language test set, in order to analyse the impact on the performance for each language. The data splits used correspond to those from monolingual experiments.
@@
Zero-shot cross-lingual: in this scenario we explore to what extent MLMs are able to generalize knowledge and metaphor transfer between these two languages in question. Therefore, we fine-tune the models with Meta4XNLI data in one language and evaluate it on the test set of the other. Data partitions used for these experiments are the same as in monolingual an multilingual scenarios.
@@
We carried out two sets of experiments to evaluate metaphor interpretation within the NLI task. We performed hyperparameter tuning, with the same range of parameters specified on the task of metaphor detection. We report best results obtained on the development set, after 4 epochs, batch size of 8, learning rate of 1e-5, weight decay of 0.1 and 512 sequence length.
The purpose of the first experiments is to examine if the presence of metaphorical expressions in premise-hypothesis pairs impacts the performance of models in the NLI task. To that end, we fine-tuned the MLMs for the task with the MultiNLI dataset. Then, we evaluated them with Meta4XNLI. Among each source dataset, we discriminated pairs with metaphors relevant to the inference relationship from those without metaphorical expressions. We developed the evaluation on each of these subsets and for each language separately, e.g. one evaluation on XNLIdev metaphor set and another on XNLIdev without metaphors set, in EN and then in ES.
The goal of the second set of experiments is to analyse the effect on models performance of not being “exposed” to instances with metaphorical expressions during training. With this aim in mind, we extracted pairs with and without metaphors from Meta4XNLI train, dev and set splits. On the first scenario, we fine-tuned the models only with pairs from the train set that did not contain any metaphorical expression. On the second scenario, we mixed pairs with and without metaphors and fine-tuned the models as well. In both cases, we evaluated on the test sets with and without metaphors for each language.
@@ 5 Results
In addition to F1 score, we performed in-vocabluary (Inv) and out-of-vocabulary (Oov) evaluations to test the impact of the labels seen from training data during the learning process of the MLMs. Precision and Recall metrics are reported in Appendix B. For in-vocabulary evaluation, we computed F1 score taking into account predicted words seen in training labeled as metaphorical. While for out-of-vocabulary evaluation we considered only words not seen during training to compute the F1 score. We included this evaluation in all experiments but for the zero-shot cross-lingual setup, since the data of train and test sets are in different languages, thus the match of the exact same metaphorical token in both partitions is highly unlikely.
Although the purpose of our experiments is not to beat state-of-the-art results but to evaluate the performance of MLMs on the task from a cross-lingual approach, we added two indicative baselines: on one hand, the system BasicBERT Li et al. (2023a), which obtained 73.3 F1 score; on the other, the result of DeBERTa reported in Sanchez-Bayona and Agerri (2022), with 73.79 F1. We selected these results for comparison purposes, since they both were evaluated on the VUAM-2020 version of the dataset in EN used in the Shared Task 2020 Leong et al. (2020) and in the same experimental setup we propose.
Cross-domain: We evaluated models trained on CoMeta and VUAM datasets with XNLI and esXNLI in ES and EN, respectively. From results reported in Table 7, we can observe that in all cases mDeBERTa outperforms XLM-RoBERTa for ES. In EN the best result is obtained by DeBERTa, which also outperforms XLM-RoBERTa in all scenarios. In all datasets, except for XNLIdev in ES, premises sentences achieve better results than hypotheses and the combination of both. This goes in line with annotation statistics in Tables 4 and 5, which show that premises contain a greater ratio of metaphors per sentence than hypotheses in both languages. In ES, in-vocabulary evaluation outperforms the general F1 score while out-of-vocabulary evaluation results decrease. The small distance in points of this cross-domain evaluation in ES exhibit stability of the models when it comes to predicting metaphors and coherence in annotations between both datasets despite the difference of text domains. Nevertheless, EN results do not demonstrate such consistency, as out-of-vocabulary obtains higher results than the overall F1 score. This might be due to the different labeling criteria used in VUAM and Meta4XNLI, as we mentioned in 3.2. These discrepancies are also reflected in a significant drop in performance with respect to ES experiments and in-domain evaluation with VUAM. The high recall scores from Table 16 in Appendix B show that the model tends to predict many metaphors, however, low precision scores indicate that a small amount of these predictions are correct.
@@
Monolingual: Results from this set of experiments are specified in Table 8. After fine-tuning and evaluating the models with Meta4XNLIES and Meta4XNLIEN, the highest overall F1 score is obtained by mDeBERTa in ES. On the other hand, XLM-RoBERTa achieves better performance than mDeBERTa in EN but still lower than in ES. In both languages, in-vocabulary evaluation results are higher than the overall ones and out-of-vocabulary results, lower. This is not the case when we use VUAM corpus for testing the model fine-tuned with Meta4XNLIEN. In this setup, similarly to results from cross-domain experiments, performance drastically falls to 32.42 F1 score and out-of-vocabulary results are the highest. The reason behind might be the mismatches in the annotation process. In ES, when evaluating CoMeta, we also encounter a decrease in the performance, however, it is more subtle: from 67.17 to 56.24, obtained by XLM-RoBERTa. This might be caused by the variety of text genres and dissimilarities between sentences from each dataset.

### v2

@@ 3.3 Resulting Dataset
The parallel data for this task comprises a total of 13320 sentences annotated at the token level, since we approached metaphor detection as a sequence labeling task, following the criteria of the cited previous work in Section 2.
With respect to Spanish annotations, there is a total of 1155 metaphorical tokens in premises and 1139 in hypotheses. Out of the 13320 sentences, 1873 contain at least one metaphorical expression, which constitutes 14% of the whole dataset. More details in Table 3 according to the various data sources and splits. The premises present a higher proportion of metaphors. This might be due to the fact that premises are longer than hypotheses (Conneau et al., 2018). In addition, the premises were collected from naturally occurring utterances, while hypotheses were generated by crowd-sourced workers. Therefore, hypothesis sentences might tend to be shorter and simpler.
@@
Regarding English annotations, we can observe a similar trend to that of ES in Table 4. A total of 3330 tokens were labeled as metaphors, and 2736 sentences contain at least one metaphorical instance out of the total 13320 sentences. Premises show a higher metaphor ratio than hypotheses, as in ES annotations. Additionally, we observe a larger amount of labeled metaphors in EN, which we attribute to the different annotation processes specified in Subsection 3.2. More specifically, given that VUAM contains a significantly higher amount of labeled metaphorical expressions, the MLM fine-tuned with this dataset predicted many ambiguous metaphors, which we partially removed in a manual revision. These discrepancies are not only noticeable in the annotations but also in the experimental results. Moreover, this gives an account of how guidelines for metaphor identification labeling are open to discussion and clarification, due to the subjective and nuanced nature of this cognitive-linguistic phenomenon.
@@
Annotations for this task were developed at the premise-hypothesis level. As shown in Table 5, the average percentage of pairs with metaphors relevant to the inference relationship is 12%. This figure remains steady throughout each source dataset and inference labels. esXNLI shows a higher number of metaphor occurrences that might be caused by the difference in text domains and sentence characteristics with respect to XNLI data. Regarding non-relevant cases, we do not use them in this work to be able to focus on establishing whether metaphor presence impacts models’ performance. We keep the same sample distribution for both languages in every experiment.
@@ 3.4 Inter-annotator Agreement
In a first round of annotation, the main annotator 1 revised the English and Spanish splits as explained in the previous section. In order to calculate inter-annotator agreement (IAA), we selected a subset of 1000 pairs in ES. This subset was labeled by annotator 2, another Spanish native speaker proficient in English and with a linguistics background. Annotator 2 labeled this subset from scratch and according to MIPVU guidelines in the case of detection, and the annotation process specified in Subsection 3.2 for interpretation. We computed Cohen’s Kappa score Cohen (1960) and obtained 0.74 in premises and 0.77 in hypotheses sentences for detection; for interpretation, the IAA score was 0.64.
We performed a second round of annotation to compare the quality of the annotations semi-automatically generated with full manual labeling. Two hired professional annotators (native Spanish with proficient knowledge of English) manually labelled 6660 instances in English and Spanish (13320 in total). This second iteration resulted in an agreement of 0.452 (Cohen’s kappa between the main original annotator 1 and the new hired annotator 4) and a Fleiss (among all three main annotators) of 0.428 (moderate agreement).
Table 6 presents the inter-annotator agreement scores of the three main annotators for metaphor detection in English and Spanish. The Cohen’s Kappa values reach a fair-moderate agreement (0.373 Cohen’s Kappa) in English, namely, between the new annotations (annotator 3) and the original annotations (semi-automatic) revised by annotator 1. Moreover, IAA between the two manual annotations (annotators 3 and 4) is quite similar, namely, 0.413 (Cohen’s Kappa, moderate agreement). Although for Spanish the IAA scores are slightly higher, they exhibit the same patterns. This shows the consistency of our original annotation with respect to the fully manual one introduced in this second iteration.
Still, metaphor annotation is inherently subjective and context-dependent, and these results reflect that complexity. The moderate kappa values obtained by the second iteration show that while annotators agree on many instances, there remain ambiguous cases that challenge consistent labeling. Furthermore, the differences between the IAA rates obtained between the two sets of annotators (original and the newly hired ones) can be attributed to the longer experience of the original two annotators in collaborating in the manual labeling of metaphor detection. Nonetheless, the obtained IAA scores show more than a moderate agreement and are comparable to those obtained in the annotation of other datasets for metaphor processing Jang et al. (2014); Kesarwani et al. (2017); Sánchez-Montero et al. (2025).
| Language | Cohen’s (1–3) | Cohen’s (1–4) | Cohen’s (3–4) | Fleiss’ |
| EN | 0.373 | 0.369 | 0.413 | 0.374 |
| ES | 0.404 | 0.452 | 0.438 | 0.428 |
3.5 Lexical Properties and Cross-Lingual Transfer
Lexical distribution
| ES | EN | |||||
| Train | Dev | Test | Train | Dev | Test | |
| Metaphor tokens | 1053 | 474 | 767 | 1527 | 697 | 1106 |
| Once-occurring metaphor tokens | 722 | 387 | 596 | 962 | 536 | 802 |
| Sentences (2 metaphors) | 126 | 80 | 116 | 198 | 110 | 161 |
| Avg sentence length (tokens) | 10.94 | 13.28 | 14.57 | 10.46 | 12.64 | 13.81 |
We conducted a corpus analysis with a focus on lexical properties and the distribution of metaphorical expressions to better understand the dataset’s characteristics. Table 7 presents the number of metaphorical tokens in the training, development, and test sets for each language, along with counts of metaphorical tokens that occur only once, the number of sentences containing multiple metaphors, and the average sentence length. The analysis reveals that English contains a higher number of metaphorical tokens overall, whereas Spanish sentences tend to be longer on average.
To better comprehend models’ performance and their generalization abilities, we examined the token overlap between training and test sets using exact match, as shown in Figures 5 and 5. In English, overlapping tokens account for 22.42% of the test set, while in Spanish, the overlap is 18.77% in the test set. These figures indicate a relatively low metaphor overlap. A more informative analysis could be conducted at the conceptual level in the future.
Examples of this overlap are illustrated in Figures 7 and 7, which display the most frequent metaphorical tokens within each partition. Many of the most frequent metaphor tokens are shared between the training and test sets, and in some cases, across both languages. These high-frequency tokens often correspond to conventional metaphorical expressions, which may influence the results of metaphor detection and interpretation tasks.
Cross-lingual transfer
We analysed metaphor transfer across languages, with quantitative results presented in Table 8. As discussed in Section 3.2, the higher number of metaphor annotations in English can be attributed to the semi-automatic annotation process applied to this language. In particular, the model used to identify metaphors in English was fine-tuned on the VUAM corpus, which contains a significantly higher proportion of labeled metaphorical tokens Sanchez-Bayona and Agerri (2022).
After manual inspection of individual cases, shown in Table 9, we found that one of the primary factors contributing to metaphor transference is the translation process. For example, instances 1, 2, 3, and 7 illustrate cases where metaphors present in the original version were expressed literally in the translation. These changes appear to be introduced during the translation process, carried out by humans in these datasets. In instance 4, for example, the metaphor “barriers” is translated as “obstáculos” in Spanish, which is a difficulty, but not a physical object that hinders advancement, like a “barrier” (“barrera” in Spanish). In contrast, instance 5 shows a case where the metaphorical expression is maintained in both languages.
The transference of some metaphorical expressions also stems from language-specific expressions or collocations. In instance 7, for example, the English verb “to spend” from the MONEY domain is used for the TIME domain. However, in Spanish, this verb does not naturally occur to express this meaning. Another language-specific phenomenon is illustrated in instance 6: while both languages use a metaphorically equivalent expression, English relies on phrasal verbs that are MWEs, whereas Spanish typically encodes the same meaning in a single token.
| ES Met | EN Met | ES ✓EN ✗ | ES ✗EN ✓ | ||
| XNLIdev | Hyp | 276 | 442 | 9 | 175 |
| Prem | 276 | 384 | 18 | 126 | |
| XNLItest | Hyp | 551 | 864 | 7 | 320 |
| Prem | 476 | 679 | 30 | 233 | |
| esXNLI | Hyp | 312 | 411 | 35 | 134 |
| Prem | 403 | 550 | 34 | 181 | |
| Train | - | 1053 | 1527 | 49 | 523 |
| Dev | - | 474 | 697 | 34 | 257 |
| Test | - | 767 | 1106 | 50 | 389 |
| # | Spanish | English | Met. Tokens (ES) | Met. Tokens (EN) | Source |
| 1 | Britney y su familia han estado desconectando en Malibú. | Britney and her family have been relaxing in Malibu. | desconectando | — | esXNLI |
| 2 | El mundo es en gran parte fruto de la ciencia. | The world is in large part a product of science. | fruto | — | esXNLI |
| 3 | La COPE y Esradio tienen una relación tormentosa. | COPE and Esradio have a difficult relationship. | tormentosa | — | esXNLI |
| 4 | Dos años después, aún existían obstáculos para compartir la información. | Information sharing barriers were still in place after two years. | — | barriers | XNLIdev |
| 5 | El director no se creía que las barreras de intercambio de información se tuvieran que eliminar completamente. | The Director did not believe that information sharing barriers should be entirely removed. | barreras | barriers | XNLIdev |
| 6 | Tienes que desarrollar cuidadosamente una organización de CIO o se derrumbará. | You have to develop a CIO organization carefully or it will fall apart. | derrumbará | fall, apart | XNLItest |
| 7 | Pasó tiempo en los Estados Unidos. | He spent time in the United States. | — | spent | XNLItest |
4 Experimental setup
In this section, we present the settings for the experiments designed to test the capabilities of multilingual MLMs (or encoder-only models) and LLMs (or decoder-only models). We evaluated metaphor detection with MLMs in cross-domain, cross-lingual, and multilingual scenarios. To assess LLMs, we additionally fine-tuned a series of models in monolingual and multilingual settings. We limited the experimental scenarios with LLMs due to the computational costs. Concerning metaphor interpretation, we also experiment with their ability to perform NLI when the correct inference requires understanding metaphorical language. We used the same encoders as for detection. In the case of decoders, we added larger models for inference with zero-shot and chain-of-thought (CoT) prompts, available in Appendix Table 29.
With the aim of developing cross-lingual experiments, we chose the multilingual language encoder-only models and checkpoints that obtained best results in Sanchez-Bayona and Agerri (2022): mDeBERTa (base) and XLM-RoBERTa (large) Conneau et al. (2020), the multilingual versions of DeBERTa He, Gao, and Chen (2021) and RoBERTa Liu et al. (2019), respectively, both for detection and interpretation. For decoder-only models, we fine-tuned Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Team (2024) and gemma-7b-it Team et al. (2024) for metaphor detection experiments, since models with a larger number of parameters are computationally expensive. Fine-tuning LLMs was based on the method by García-Ferrero et al. (2024), which enables fine-tuning and inference on LLMs for sequence labeling tasks with constrained decoding. In the case of metaphor interpretation, we evaluated Llama-3.3-70B-Instruct, Qwen2.5-72B-Instruct, gpt-4o through in-context learning in order to test LLMs understanding capabilities. Experiments with open-weight models were performed via the HuggingFace implementations. For gpt-4o 222https://platform.openai.com/ we used the API.
@@ 4 Experimental setup
Taking advantage of previously available resources and Meta4XNLI, we conducted a series of experiments to evaluate and fine-tune MLMs and LLMs. The configuration of cross-domain experiments is specified in Sanchez-Bayona and Agerri (2022). For the other three setups with encoders, monolingual, multilingual, and zero-shot cross-lingual experiments, we performed hyperparameter tuning through grid search with batch size [8, 16, 32], linear decay [0.1, 0.01], learning rate (in the [1e-5-5e-5] interval), sequence length of 128 and epochs from 4 to 10. A warm-up of 6% is specified. The results of the hyperparameter tuning showed that after 4 epochs, development loss started to increase, so the results reported here are obtained with 4 epochs, batch size of 8, weight decay of 0.1 and learning rate of 5e-5.
For the fine-tuning of decoders in mono- and multilingual setups, we also performed grid search over the development set and the following parameter space: learning rate [2e-4, 2e-5, 2e-6, 2e-7], max sequence length 512, batch size [8, 16, 32], epochs [5, 10, 30] and a fixed seed. We report the best results obtained with 30 epochs, 8 batch size and 2e-5 learning rate.
@@
Cross-domain: the aim of this set of experiments is to evaluate the performance of MLMs fine-tuned with CoMeta and VUAM datasets on Meta4XNLIES and Meta4XNLIEN, respectively, since each dataset contains texts from different domains. Specifically, text sources from XNLI cover Face-To-Face, Telephone, Government, 9/11, Letters, Oxford University Press, Slate, Verbatim, and Government and Fiction. esXNLI is a compilation of texts from 5 sources: a newspaper, an economic forum, a celebrity magazine, a literature blog, and a consumer magazine. The motivation is to explore the impact of text features and genres on the performance, as well as annotation criteria Aghazadeh, Fayyaz, and Yaghoobzadeh (2022); Lai, Toral, and Nissim (2023). To do so, we chose the models with the best performance from monolingual experiments developed in Sanchez-Bayona and Agerri (2022). We conducted the evaluation on various data splits: within each source dataset, XNLIdev, XNLItest and esXNLI, we evaluated premises and hypotheses separated and combined. This is due to the dissimilarities and the unequal distribution of metaphorical expressions between premises and hypothesis sentences mentioned in Subsection 3.3.
@@
Monolingual: this scenario comprises the fine-tuning and evaluation of MLMs and LLMs on Meta4XNLIES and Meta4XNLIEN separately. To accomplish this task, we split Meta4XNLI into train, development, and test sets (0.6-0.2-0.2). We equally distributed the data to ensure each partition is balanced in terms of source datasets, sentences, and metaphor occurrence. Data statistics are detailed in the Appendix. These partitions will be used as well in subsequent multilingual and zero-shot cross-lingual scenarios. In addition to fine-tuning and evaluation on Meta4XNLIES and Meta4XNLIEN, we evaluated each trained monolingual model with the test sets of CoMeta and VUAM, following the same reasoning as in cross-domain experiments.
@@
Multilingual: The purpose of these experiments is to explore whether MLMs and LLMs benefit from being trained on data in multiple languages. In this case, we combined Meta4XNLIES and Meta4XNLIEN train splits to fine-tune the models. Subsequently, we evaluated the trained models on each language test set, in order to analyse the impact on the performance for each language. The data splits used correspond to those from monolingual experiments.
@@
Zero-shot cross-lingual: in this scenario, we explore to what extent MLMs are able to generalize knowledge and metaphor transfer between these two languages in question. Therefore, we fine-tune the models with Meta4XNLI data in one language and evaluate them on the test set of the other. Data partitions used for these experiments are the same as in monolingual and multilingual scenarios.
@@
We carried out two sets of experiments to evaluate metaphor interpretation within the NLI task with encoder-only models. We performed hyperparameter tuning through grid search, with the same range of parameters specified on the task of metaphor detection. We report the best results obtained on the development set, after 4 epochs, batch size of 8, learning rate of 1e-5, weight decay of 0.1 and 512 sequence length.
To assess decoder-only models’ performance on the NLI task with sentences with and without metaphors, we performed inference over the specified models with the following parameters and Hugging Face and vLLM Kwon et al. (2023) implementations: max new tokens=5, temperature=0.3. Models were prompted to answer with one of the three NLI relations [entailment, natural, contradiction]. To do so, we designed two prompts available in Appendix Table 29: one with no examples (zero-shot) and another one with longer context and one example for each label (chain-of-thought (CoT)). We only report the results for the first set of experiments, due to the computational cost of fine-tuning models of this size.
The purpose of the first experiments is to examine whether the presence of metaphorical expressions in premise-hypothesis pairs impacts the performance of models in the NLI task. To that end, we fine-tuned the MLMs for the task with the MultiNLI dataset. Then, we evaluated them with Meta4XNLI. Among each source datasets, we discriminated pairs with metaphors relevant to the inference relationship from those without metaphorical expressions. We developed the evaluation on each of these subsets and for each language separately, e.g., one evaluation on XNLIdev metaphor set and another on XNLIdev without metaphors set, in EN and then in ES. The goal of the second set of experiments is to analyse the effect on models’ performance of not being “exposed” to instances with metaphorical expressions during training. With this aim in mind, we extracted pairs with and without metaphors from Meta4XNLI train, dev, and set splits. In the first scenario, we fine-tuned the models only with pairs from the train set that did not contain any metaphorical expressions. In the second scenario, we mixed pairs with and without metaphors and fine-tuned the models as well. In both cases, we evaluated on the test sets with and without metaphors for each language.
@@ 5 Results
In addition to F1 score, we computed in-vocabulary (Inv) and out-of-vocabulary (Oov) F1 scores to assess the impact of the labels seen from training data during the learning process of the models. Precision and Recall metrics are reported in Appendix C. For the in-vocabulary evaluation, we calculated the F1 score using only the predicted metaphorical tokens that also appeared in the training data labeled as metaphors. In contrast, the out-of-vocabulary evaluation was computed on predicted metaphorical tokens that were not labeled as metaphors in the training set. We used exact match to extract the overlapping tokens. We included this evaluation in all experiments, but for the zero-shot cross-lingual setup, since the data of the train and test sets are in different languages, the match of the same metaphorical token in both partitions is highly unlikely.
Although the purpose of our experiments is not to beat state-of-the-art results but to evaluate the performance of MLMs and LLMs on the task from a cross-lingual approach, we added two relevant baselines: on one hand, the system BasicBERT Li et al. (2023a), which obtained 73.3 F1 score; on the other, the result of DeBERTa reported in Sanchez-Bayona and Agerri (2022), with 73.79 F1. We selected these results for comparison purposes, since they both were evaluated on the VUAM-2020 version of the dataset in EN used in the Shared Task 2020 Leong et al. (2020) and in the same experimental setup we propose.
@@
Cross-domain: We evaluated models trained on CoMeta and VUAM datasets with XNLI and esXNLI in ES and EN, respectively. From results reported in Table 10, we can observe that in all cases mDeBERTa outperforms XLM-RoBERTa for ES. In EN the best result is obtained by DeBERTa, which also outperforms XLM-RoBERTa in all scenarios. In all datasets, except for XNLIdev in ES, premise sentences achieve better results than hypotheses and the combination of both. This goes in line with annotation statistics in Tables 3 and 4, which show that premises contain a greater ratio of metaphors per sentence than hypotheses in both languages. In ES, in-vocabulary evaluation outperforms the general F1 score while out-of-vocabulary evaluation results decrease. The small distance in points of this cross-domain evaluation in ES exhibits stability of the models when it comes to predicting metaphors and coherence in annotations between both datasets, despite the difference of text domains.
Nevertheless, EN results do not demonstrate such consistency, as out-of-vocabulary obtains higher results than the overall F1 score. These discrepancies are also reflected in a significant drop in performance with respect to ES experiments and in-domain evaluation with VUAM. Thus, the high recall scores from the Appendix Table 24 show that the model tends to predict many metaphors; however, low precision scores indicate that a small amount of these predictions are correct. Previous work showed that even though MIPVU guidelines (used to annotate VUAM) were applied for the manual annotation of the CoMeta dataset, the guidelines allow for a level of subjectivity in their interpretation, which means that the proportion of metaphorical tokens was around 6% of the total for VUAM, while in CoMeta it is around 2% Sanchez-Bayona and Agerri (2022).
Regarding Meta4XNLI, the Spanish data was fully revised manually, while, for English, we used a semi-automatic approach (described in Section 3.2) based on first deploying encoders fine-tuned on English VUAM data and then a manual revision of the automatic annotations. This means that, during the manual revision phase, the English dataset already contained preliminary labels automatically annotated with models fine-tuned on VUAM. As a consequence, when comparing the English with the manually annotated Spanish split, and just as it happened between CoMeta and VUAM, the English split contains a higher number of annotated metaphorical tokens. Thus, while the differences are not as large as in the case of CoMeta and VUAM (where domain differences need to be factored in), the English Meta4XNLI contains around 6% more metaphorical tokens than the Spanish split.
Monolingual: Results from this set of experiments are specified in Tables 11 and 12. After fine-tuning and evaluating the models with Meta4XNLIES and Meta4XNLIEN, the highest overall F1 score is obtained by mDeBERTa in ES for encoders, and Llama-3.1-8B-Instruct for decoders. On the other hand, XLM-RoBERTa and gemma-7b-it achieve the best performance in EN but still lower than in ES.